OpenStudy (pawanyadav):

The solution of differential equation.. (x^2-y^2)dx +2xy dy=0

OpenStudy (samigupta8):

Is the ans 1/2 ( y/x)^2=ln(c/x)

OpenStudy (xapproachesinfinity):

that's and exact equation

OpenStudy (loser66):

No, it is not,kid

OpenStudy (xapproachesinfinity):

why nit?

OpenStudy (loser66):

\(\partial M/\partial y=-2y\neq \partial N/\partial x= 2y\)

OpenStudy (pawanyadav):

Yes , that equation is not right

OpenStudy (loser66):

So, you have to find \(\mu (x)\) or \(\mu (y)\) or \(\mu(xy)\) to have is exact I worked on it, \(\mu (x)=1/x^2\) works well

OpenStudy (xapproachesinfinity):

yes, i miss that minus

OpenStudy (samigupta8):

Sorry ! I didn't take that 2 into consideration..

OpenStudy (loser66):

@Pawanyadav use (1/x^2) to multiple into both sides and then solve as usual

OpenStudy (pawanyadav):

I do the same but can't reach the answer

OpenStudy (loser66):

Show your work, please.

OpenStudy (samigupta8):

I got this y^2+x^2=cx

OpenStudy (xapproachesinfinity):

yeah old man is right

OpenStudy (pawanyadav):

I got logx= log (1+ (y/x)^2)+c Whats next

OpenStudy (xapproachesinfinity):

so what is the problem here, you made it exact

OpenStudy (samigupta8):

Not log x .. It has to be log (1/x)

OpenStudy (xapproachesinfinity):

and solved it?

OpenStudy (loser66):

kid, then solve for f(x,y)

OpenStudy (xapproachesinfinity):

yes i know old man, I was talking to the poster

OpenStudy (loser66):

kid, take over, please :)

OpenStudy (loser66):

@samigupta8 you got it right

OpenStudy (xapproachesinfinity):

\[(1-\frac{y^2}{x^2})dx+2\frac{y^2}{x}dy=0\]

OpenStudy (xapproachesinfinity):

which one did you integrate

OpenStudy (pawanyadav):

@samigupta how its log (1/x)

OpenStudy (xapproachesinfinity):

that supposed to have only y not y^2

OpenStudy (samigupta8):

I got these equations v/2 + xdv/dx +1/2v=0 So now u take all v terms on left side and x terms on right.. Solve it u will get the desired ans

OpenStudy (samigupta8):

v is y/x here

OpenStudy (xapproachesinfinity):

\[f(x,y)=\int 1-\frac{y^2}{x^2}dx=x+y^2/x+h(y)=c\]

OpenStudy (loser66):

Then, take derivative of the above result w.r.t y it is = 2y/x +h'(y)

OpenStudy (xapproachesinfinity):

yeah i know, just looking for something

OpenStudy (loser66):

Ok, kid, please, give him step by step. (Whole thing, please).

OpenStudy (pawanyadav):

On dividing by x^2 dx,we got (1-(y/x)^2)=-2(y/x)y' Now put y=tx I got (1-t^2)/-2t=t+x(dt/dx)

OpenStudy (pawanyadav):

Then (t^2-1)/2t -t=x(dt/dx)

OpenStudy (xapproachesinfinity):

one second i will write the procedure

OpenStudy (pawanyadav):

Finally dx/x=(-2t/(1+t^2))dt On integrate -logx =log (1+t^2)+c

OpenStudy (pawanyadav):

Thanks ,, finally I got it

OpenStudy (pawanyadav):

Thanks old man and kid ..your method is little difficult..

OpenStudy (xapproachesinfinity):

what method are you using ?

OpenStudy (pawanyadav):

Its all above..

OpenStudy (xapproachesinfinity):

yeah don't get, just tell me the name, i will check on it

OpenStudy (pawanyadav):

Homogenisation of equation

OpenStudy (xapproachesinfinity):

oh i see

OpenStudy (pawanyadav):

First convert to homogeneous equation than solve by variable separable form

OpenStudy (xapproachesinfinity):

so you get the solution i wrote

OpenStudy (xapproachesinfinity):

yeah i have seen that before just don;t recall the whole procedure now

OpenStudy (pawanyadav):

No ,I can't open this page

OpenStudy (xapproachesinfinity):

it a word document

OpenStudy (pawanyadav):

OK ..I'll get i

OpenStudy (xapproachesinfinity):

i made it a PDF

OpenStudy (pawanyadav):

Thanks , now I can opeopen

OpenStudy (pawanyadav):

I never solve any questions like this before

OpenStudy (loser66):

bravo, kid